### v1

# Nonlocal Locking of Observable Quantities: A Faithful Signature of Nonclassical Correlations

Abstract
Nonclassicality in composite quantum systems depicts several puzzling manifestations, with Einstein-Podolsky-Rosen entanglement, Schrödinger steering, and Bell nonlocality being the most celebrated ones. In addition to those, an unentangled quantum state can also exhibit nonclassicality, as evidenced from notions such as quantum discord and work deficit. Here, we propose a general framework to investigate nonclassical correlations in multipartite quantum states. The distinct signatures left on observable quantities, depending on whether the sub-parts of a composite system are probed separately or jointly, provide an operational avenue to construct different quantifiers that faithfully capture signatures of nonclassicality in quantum states. Along the line we unveil an intriguing phenomenon referred to as ‘nonlocal locking of observable quantities’, where the value of an observable quantity gets locked in the correlation of a nonclassical state. Our approach reduces the experimental demand for verification of nonclassicality in composite systems and can find applications for enhaced energy storage in quantum thermodynamical devices.
Introduction.– Correlations stand at the forefront of scientific inquiry across various fields of study. Their systematic investigation plays a crucial role in enhancing our understanding of the physical world [1]. This is especially evident while studying quantum systems consisting of multiple sub-parts. For instance, the puzzling features of quantum correlations as exemplified by the phenomena of entanglement and steering have been recognized since the early days of quantum theory [2, 3, 4, *Schrdinger1936]. Later J. S. Bell, in his seminal work, built on the notion of quantum entanglement to establish the physical phenomenon of quantum nonlocality [6, *Bell1966] (see also [8]). Along with capturing fundamentally distinct notions of nonclassicality [9, 10, 11], quantum entanglement, steering, and nonlocality also find a number of utilities across a spectrum of applications at the heart of modern quantum technologies [12, 13, 14, 15]. More recently, it has been identified that quantum discord encapsulates nonclassicality even in an unentangled quantum state [16, 17, 18] (see also [19, 20]). The study of nonclassical correlations, with an emphasis on their identification, quantification, and classification, thus stands as a pivotal objective in quantum information research during the last few decades.
In this paper, we adopt an operationally-driven approach to explore nonclassical correlations in multipartite quantum systems. A multipartite quantum system can be probed in different ways depending on the availability of resources among the spatially separated sub-parts. For instance, global probing requires quantum communication channels among the sub-parts, and without such channels only local probing is possible, where distant parties address their respective sub-parts individually. On the other hand, allowing only classical communication among the sub-parts confines a probing mechanism intermediate to the global and local probings (see Fig. 1). A correlated quantum state can lead to distinct observable values depending on how the system is probed and this observation provides operational pathways for exploring nonclassical correlations within it. In this work, we pursue this approach to introduce quantifiers of general nonclassical correlations beyond entanglement. While some quantifiers are derived from resource theories and hinge on entropic quantities, the generality of our framework allows for constructing alternative quantifiers leveraging simple observables on composite systems, thus mitigating experimental complexity. For bipartite systems, such quantifiers are proven to be faithful, meaning that they vanish on and only on states lacking nonclassical correlations. Our analysis unveils an intriguing phenomenon hereby termed ‘nonlocal locking of observable quantities,’ wherein observable values get locked within the nonclassical correlation inherent in multipartite quantum states. Furthermore, we provide exact quantitative results focusing on the simplest composite quantum system comprising two qubits. Finally, we discuss on the potential implications of our findings in designing efficient energy storage devices that utilize nonclassically correlated states of many-body quantum systems.
Preliminaries.– The state of an individual quantum system ($A$) is described by a density operator $\rho\in\mathcal{D}(\mathcal{H}_{A})$; $\mathcal{D}(\mathcal{X})$ be the set of positive trace-one operators acting on the Hilbert space $\mathcal{X}$. For a finite-dimensional system, which we assume throughout this work, $\mathcal{H}_{A}$ is isomorphic to some complex Euclidean space $\mathbb{C}^{d_{A}}$, where $d_{A}$ denotes the dimension of the Hilbert space. The state of a composite system ${\bf A}\equiv A_{1}\cdots A_{n}$ is called separable if it is a statistical mixture of independently prepared constituent sub-parts, i.e., $\rho_{\bf A}=\sum_{i}p_{i}(\otimes_{j=1}^{n}\psi^{i}_{j})$, where $\vec{p}:=\{p_{i}\}$ is a probability vector, and $\psi:=\ket{\psi}\bra{\psi}$ | with |$\ket{\psi^{i}}_{j}\in\mathbb{C}^{d_{j}}$. States that are not separable are called entangled, and they capture the most celebrated aspect of nonclassical correlation in composite quantum systems [12, 13]. Importantly, a separable state can also accommodate another distinct nonclassical aspect of correlation recognized as quantum discord [16, 17, 18]. A multipartite state is called classically correlated if it allows a decomposition of the form $\rho_{\bf A}=\sum p_{i_{1}\cdots i_{n}}\psi^{i_{1}}_{1}\otimes\cdots\otimes\psi^{i_{n}}_{n}$, where {|$\{\ket{\psi^{i_{j}}}_{j}\}$ forms an orthonormal basis (ONB) for the $j^{th}$ Hilbert space, and $\{p_{i_{1}\cdots i_{n}}\}$ denotes a probability vector. A function $\mathbb{F}:\mathcal{D}(\otimes_{j=1}^{n}\mathbb{C}^{d_{j}}_{A_{j}})\mapsto\mathbb{R}$ is said to capture the signature of nonclassical correlation faithfully if it takes value zero for all classically correlated states and nonzero otherwise. For the bipartite case, the original entropic measure of quantum discord, defined in terms of the difference between two different expressions for the mutual information [16, 17, 18], is an example of a faithful quantifier [20]. The notion of quantum discord has been generalized to quantum systems involving more than two sub-parts, albeit with caveats in its operational interpretation [21, 22, 23]. Notably, this notion of nonclassicality is party asymmetric, as a state can be a statistical mixture of product states where {|$\{\ket{\psi^{i_{j}}}\}$ form ONBs only for some sub-parts, but not for all. In the following, we develop a general operational framework for studying nonclassical correlations in multipartite quantum states.
Framework.– Any physical operation acting on a quantum system is described by a completely positive trace-preserving (CPTP) map [24]. Given a quantum system $A$, we start by defining a generic function $f_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathbb{R}$. For a CPTP map $\phi_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathcal{D}(\mathbb{C}^{d_{A}})$, we define the variation of the function $f_{A}$ on some state $\rho_{A}$ as $\Delta f_{A}^{\phi_{A}}(\rho_{A}):=f_{A}(\rho_{A})-f_{A}(\phi(\rho_{A}))$. For a set of CPTP maps $\Phi_{A}$, the variation is given by $\Delta f_{A}^{\Phi_{A}}(\rho_{A}):=\max_{\phi_{A}\in\Phi_{A}}\Delta f_{A}^{\phi_{A}}(\rho_{A})$, which quantifies the maximum extracted value of the function $f_{A}$ on the given state $\rho_{A}$ under the set of operations $\Phi_{A}$. For a multipartite quantum system ${\bf A}(=A_{1}\cdots A_{n})$, fixing $(f_{j},\Phi_{j})$ pair for the $j^{th}$ party $j\in\{1,\cdots,n\}\equiv[n]$, and fixing $(f_{\bf A},\Phi_{\bf A})$ pair for the joint system, we define another function $\mathbb{F}_{GL}(\rho_{\bf A}):=\Delta f^{\Phi_{\bf A}}_{\bf A}(\rho_{\bf A})-\sum_{j}\Delta f^{\Phi_{j}}_{j}(\rho_{j})$, where $\rho_{j}:=\operatorname{Tr}_{A_{j}^{c}}(\rho_{\bf A})$ denotes the marginal state of the $j^{th}$ sub-part. The function $\mathbb{F}_{GL}$ quantifies the difference between the extracted value of the global function and that of the local functions evaluated on the global state and the local sub-parts under the fixed sets of operations. Interestingly, as we will see, for suitable choices of $(f_{\gamma},\Phi_{\gamma})$’s, with $\gamma\in\{[n]\cup{\bf A}\}$, the function $\mathbb{F}_{GL}$ can capture the signature of different forms of correlation within the multipartite state. As an example, for bipartite systems, the following observation is noteworthy (see Appendix-I for more details).
Figure 1: Resource Extraction from a bipartite quantum state $\rho_{\bf A}\in\mathcal{D}(\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}})$; here ${\bf A}=A_{1}A_{2}$. (a) $\Delta f^{\Phi_{12}}_{12}(\rho_{\bf A})$ quantifies the amount of resource extracted under the set of global operations $\Phi_{12}$. (b) $\sum_{j}\Delta f^{\Phi_{j}}_{j}(\rho_{j})$ is the amount of resource extracted under local operations $(\Phi_{1},\Phi_{2})$. (c) $\Delta f^{\{\Phi_{\gamma}\}}_{\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ is the amount of resource extracted under $(\Phi_{1},\Phi_{2},\Phi_{12})$ when from the sub-part $A_{1}$ to $A_{2}$ free classical channel $\mathcal{N}^{0}_{C}$ are available. The quantity $\mathbb{F}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A}):=\Delta f^{\Phi_{12}}_{12}(\rho_{\bf A})-\Delta f^{\{\Phi_{\gamma}\}}_{\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ amounts to ’nonlocal locking of the resource’, and can capture the signature of nonclassical correlation in state $\rho_{\bf A}$.
Observation 1.
The choices $(f_{\gamma},\Phi_{\gamma})\equiv(f_{\gamma}^{pe},\Phi_{\gamma}^{no}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \gamma\in\{[2]\cup{\bf A}\}$ lead to $\mathbb{F}^{pe}_{GL}(\rho_{\bf A})=S(\rho_{1})+S(\rho_{2})-S(\rho_{\bf A}):=I(A_{1}:A_{2})$; where $S(\rho_{\gamma}):=-\mathrm{tr}(\rho_{\gamma}\log\rho_{\gamma})$, $f^{pe}_{\gamma}(\rho_{\gamma}):=\log d_{\gamma}-S(\rho_{\gamma})$, and $\Phi^{no}_{\gamma}$ is the set of noisy operations.
Notably, the function $f^{pe}_{\gamma}(\rho_{\gamma})$ denotes the asymptotic rate of purity extraction from the state $\rho_{\gamma}$ under noisy operations [25]. Accordingly, $\mathbb{F}^{pe}_{GL}$ in Observation 1, quantifies the difference between asymptotic rates of purity extraction under global and local noisy operations [26], which interestingly finds an exact match with the total correlation of the state $\rho_{\bf A}$, quantified through the quantum mutual information $I(A_{1}:A_{2})$ [27].
At this point it is worth noting how the function $\mathbb{F}_{GL}$ be evaluated operationally for a multipartite state (see Fig. 1). Firstly, each of the local parties can evaluate the corresponding $\Delta f^{\Phi_{j}}_{j}$ locally. On the other hand, for evaluating $\Delta f^{\Phi_{\bf A}}_{\bf A}$, the spatially separated parties can send their respective sub-parts to a common party (say $j^{\star}$). This however demands a perfect quantum communication channel from all the other parties to the designated party $j^{\star}$. At this point, one can consider intermediate scenarios where only a subset of channels $\mathcal{N}$ are available from the remaining parties to $j^{\star}$. In such cases one can define the variation $\Delta f^{\{\Phi_{\gamma}\}}_{\mathcal{N}}(\rho_{\bf A}):=\max_{\{\phi_{\gamma}\in\Phi_{\gamma},\leavevmode\nobreak\ N\in\mathcal{N}\}}[\sum_{j\in[n]}\Delta f^{\phi_{j}}_{j}(\rho_{j})+\Delta f^{\phi_{\bf A}}_{\bf A}(N(\sigma_{\bf A}))-\mathcal{C}(N)]$, where $\sigma_{\bf A}:=(\otimes_{j}\phi_{j})(\rho_{\bf A})$, $N(\sigma_{\bf A})$ is the composite state received by $j^{\star}$, and $\mathcal{C}(N)$ is a suitably defined cost function. This function quantifies the value extraction from a given $\rho_{\bf A}$ over the maps $\{\Phi_{\bf A}\}$’s when collaboration from the remaining parties to $j^{\star}$ is limited by the set of maps $\mathcal{N}$. To say more elaborately, each of the parties extracts $\Delta f^{\phi_{j}}_{j}$ from their respective parts under their local operations resulting the state $\rho_{\bf A}$ to evolve into a new state $\sigma_{\bf A}$. Accordingly, the $j^{\star}$ party receives the state $N(\sigma_{\bf A})$ and extracts $\Delta f^{\phi_{\bf A}}_{\bf A}$. Clearly, in absence of $\mathcal{N}$, the value of $\Delta f^{\{\Phi_{\gamma}\}}_{\mathcal{N}}$ should exactly match with $(\sum_{j\in[n]}\Delta f^{\Phi_{j}}_{j})$, whereas with perfect quantum communication channel it should match with $\Delta f^{\Phi_{\bf A}}_{\bf A}$. The cost function $\mathcal{C}(N)$ is defined accordingly to satisfy these demands [28, 29]. At this point, we define a new function $\mathbb{F}_{G\mathcal{N}}(\rho_{\bf A}):=\Delta f^{\Phi_{\bf A}}_{\bf A}(\rho_{\bf A})-\Delta f^{\{\Phi_{\gamma}\}}_{\mathcal{N}}(\rho_{\bf A})$ quantifying the difference between optimal global extracted value and the $\mathcal{N}$ channel assisted optimal extracted value, which will be shown to be useful to explore nonclassical correlation in multipartite states.
Resource theoretic quantifiers.– To see application of the aforesaid general framework we start by analyzing how the function $\mathbb{F}_{G\mathcal{N}}$ can capture the notion of quantum discord in bipartite systems (${\bf A}=A_{1}A_{2}$) for suitable choices of the tuple $(f_{\gamma},\Phi_{\gamma},\mathcal{N})$. For that once again we fix $(f_{\gamma},\Phi_{\gamma})\equiv(f^{pe}_{\gamma},\Phi_{\gamma}^{no}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \gamma\in\{1,2,{\bf A}\}$, and fix $\mathcal{N}$ to be classical communication channel between $A_{1}$ and $A_{2}$. Recall that quantum discord is a party asymmetric concept: a bipartite state $\rho_{\bf A}\in\mathcal{D}(\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}})$ having the classical-quantum (CQ) form $\rho_{\bf A}=\sum p_{i_{1}i_{2}}\psi^{i_{1}}_{1}\otimes\psi^{i_{2}}_{2}$, with $\langle\psi^{i_{1}}_{1}|\psi^{i^{\prime}_{1}}_{1}\rangle=\delta_{i_{i}i^{\prime}_{1}}$, has discord zero from $A_{1}\to A_{2}$ [16, 17, 18]. A generic classical communication channel $N\in\mathcal{N}_{C}$ sending the $A_{1}$ part to the $A_{2}$ can be thought as a measure-and-prepare channel [30, 31], with preparations being restricted to rank-1 orthogonal states (see Appendix III). Such a channel could map completely mixed state to a resourceful state, and therefore is costly from the perspective of purity resource theory. One can bypass this cost association by further limiting these channels to be cost free. As we argue in Appendix III, a cost-free classical channel in this case takes the form $N(\rho_{A_{1}}):=\sum_{l\in[d_{1}]}\operatorname{Tr}(\pi_{l}\rho_{1})\ket{\alpha^{l}_{1}}\bra{\alpha^{l}_{1}}$ |, where $\langle\alpha^{l}_{1}|\alpha^{l^{\prime}}_{1}\rangle=\delta_{ll^{\prime}}$, and $\{\pi_{l}\}$ is a positive-operator-value-measurement (POVM) with $\operatorname{Tr}(\pi_{l})=1,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ l\in[d_{1}]$. By denoting the set of all such zero cost classical channels as $\mathcal{N}^{0}_{C}$, we obtain the following result.
Theorem 1.
For $(f_{\gamma},\Phi_{\gamma},\mathcal{N})\equiv(f_{\gamma}^{pe},\Phi_{\gamma}^{no},\mathcal{N}^{0}_{C})$ the function $\mathbb{F}^{pe}_{G\mathcal{N}^{0}_{C}}$ faithfully captures nonclassical correlation from $A_{1}\to A_{2}$ in any bipartite state $\rho_{\bf A}\in\mathcal{D}(\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}})$ having maximally mixed marginal on $A_{1}$ sub-part.
Proof is provided in Appendix III. During the process we also prove the following Proposition which is interesting on its own right (see Appendix II).
Proposition 1.
For $\Lambda:\mathcal{D}(\mathbb{C}^{d})\to\mathcal{D}(\mathbb{C}^{d})$ being a CPTP unital map, the states $\rho$ and $\Lambda(\rho)$ have the same spectrum whenever $S(\Lambda(\rho))=S(\rho)$.
At this point we would like to point out that a special case of our Theorem 1 reproduces one of Zurek’s results [32]. While the set of free channels there consists of measure-and-orthogonal preparation with the measurement being rank-1 projective measurements, in our case we consider more general kind of measurements. Notably, similar analysis as of Theorem 1 is worth exploring by fixing $(f_{\gamma},\Phi_{\gamma},\mathcal{N})\equiv(f_{\gamma}^{fe},\Phi_{\gamma}^{to},\mathcal{N}_{tc})$, where $f_{\gamma}^{fe}$and $\Phi_{\gamma}^{to}$ respectively denote thermodynamic free energy and thermal operations for a given inverse temperature $\beta$, and $\mathcal{N}_{tc}$ are thermal classical channels compatible to this thermodynamic resource theory. The resulting study will go in line with the works in Refs. [33, 34, 35, 36].
Generic nonclassicality quantifiers.– We now proceed to explore the generality of our framework to construct quantifiers for nonclassical correlations that need not to be motivated from a resource theory. To this aim, we start by noting that the function $f:\mathcal{D}(\mathbb{C}^{d})\mapsto\mathbb{R}$ can simply be the expectation value of some observable $\mathcal{O}\in\mbox{Herm}(\mathbb{C}^{d})$, i.e., $\mathcal{O}:\rho\mapsto\langle\mathcal{O}\rangle_{\rho}:=\operatorname{Tr}(\mathcal{O}\rho)$. Unlike the earlier case, here the variation is evaluated over the set of reversible operations, resulting $\Phi$ to be the set of unitary operations $\mathrm{U}$. In this case the variation $\Delta\mathcal{O}^{\mathrm{U}}(\rho):=\max_{u\in\mathrm{U}}[\operatorname{Tr}(\mathcal{O}\rho)-\operatorname{Tr}(\mathcal{O}u\rho u^{\dagger})]$ quantifies the optimal value of the observable $\mathcal{O}$ extracted from the state $\rho$ when the system is evolved under closed system dynamics. While studying multipartite systems, generality of our framework allows choosing different observables for different sub-parts. Although the choice of the global observable can also be arbitrary, here we consider $\mathcal{O}_{\bf A}\equiv\sum_{j\in[n]}\tilde{\mathcal{O}}_{j}\leavevmode\nobreak$, where $\tilde{\mathcal{O}}_{j}:=\mathcal{O}_{j}\otimes\mathbf{I}_{rest}$. If system’s Hamiltonian $h$ is taken to be the observable, the variation $\Delta h^{\mathrm{U}}(\rho)$, known as the ergotropy of the state, which has a long history in literature [37, 38, 39]. The corresponding function $\mathbb{H}_{GL}(\rho_{\bf A})$ quantifies the ergotropic gap – the difference between the extractable work from a state $\rho_{\bf A}$ under global and local unitary operations, respectively. Interestingly, this function and its variants can capture different forms of nonclassical correlations, particularly entanglement in bipartite states and genuine entanglement in multipartite quantum states [40, 41, 42, 43, 44].
Figure 2: Nonlocal energy locking in two qubit Werner states $\mathcal{W}^{\alpha}_{\bf A}$. Red dashed curve denotes the original quantum discord $\mathcal{Q}(\mathcal{W}^{\alpha}_{\bf A})=\frac{1-\alpha}{4}\log(1-\alpha)-\frac{1+\alpha}{2}\log(1+\alpha)+\frac{1+3\alpha}{4}\log(1+3\alpha)$ establishing the states are nonclassical for $\alpha\in(0,1]$ [17] (see also [45]). Blue curve amounts to energy $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\mathcal{W}^{\alpha}_{\bf A})\leavevmode\nobreak\ [=\alpha\epsilon_{2}/2]$ locked into the nonclassical correlation of the state $\mathcal{W}^{\alpha}_{\bf A}$. Here the plot is given by fixing $\epsilon_{2}=2$.
While studying the quantity $\mathbb{O}_{G\mathcal{N}^{0}_{C}}$, which is one of our core interests in the present work, we first need to identify the free classical channels $\mathcal{N}^{0}_{C}$. As it can be argued a classical channel $N\in\mathcal{N}_{C}$ communicating $A_{j}$ sub-part to the $j^{\star}$ party will be considered to be free if $\operatorname{Tr}(\mathcal{O}_{j}N(\rho_{j}))\leq\operatorname{Tr}(\mathcal{O}_{j}\rho_{j})\leavevmode\nobreak\ \&\leavevmode\nobreak\ \Delta\mathcal{O}^{U_{j}}_{j}N(\rho_{j})\leq\Delta\mathcal{O}^{U_{j}}_{j}(\rho_{j}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \rho_{j}\in\mathcal{D}(\mathbb{C}^{d_{j}})$. A more elaborate justification of this demand as well as a complete characterization of free classical channels $\mathcal{N}^{0}_{C}$ are provided in Appendix IV. With this we now present a general result for bipartite systems (proof is deferred to Appendix IV).
Theorem 2.
The function $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ faithfully captures nonclassical correlation from $A_{1}\to A_{2}$ in any bipartite state $\rho_{\bf A}$, whenever $\mathcal{O}_{\bf A}$ has non degenerate spectrum.
Notably unlike Theorem 1, the function $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ faithfully captures signature of nonclassical correlation in any bipartite state without assuming any further constraint on the state. Theorem 2 has quite interesting implication. A non zero value of $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ implies that extractable value of the observable $\mathcal{O}_{\bf A}$ under global operations is strictly greater than the value extractable under local operations assisted with classical communication from $A_{1}$ to $A_{2}$. In other words, the value of the observable $\mathcal{O}_{\bf A}$ gets locked in the nonclassical correlation of the state. This phenomenon we call the nonlocal locking of observable quantities in nonclassical correlation. The motivation of this name arises from the pioneering ‘nonlocality without entanglement’ phenomenon, where distinguishability of a set of multipartite product states under local operation and classical communication is suboptimal than the global distinguishability [46, 47] (see also [48, 49, 50, 51]), and the works [52, 53, 54] that report several locking phenomena utilizing quantum resources). The following crucial property (proof provided in Appendix V) of the quantity discussed in Theorem 2 is worth noting.
Proposition 2.
$\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})=\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\sigma_{\bf A})$, where $\sigma_{\bf A}=u_{1}\otimes u_{2}(\rho_{\bf A})u^{\dagger}_{1}\otimes u^{\dagger}_{2}$ for arbitrary $u_{i}$’s on the respective parts.
A simplified from of $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ can be obtained for generic two-qubit state $\rho_{\bf A}=\frac{1}{4}(\mathbf{I}+\bm{\sigma}^{\intercal}_{1}{\bf r}_{1}+{\bf r}^{\intercal}_{2}\bm{\sigma}_{2}+\bm{\sigma}^{\intercal}_{1}{\bf T}\bm{\sigma}_{2})$, which is stated in the following theorem by considering $\mathbb{O}$ to be the Hamiltonian of the system.
Theorem 3.
$\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})=\min_{\bf m}\sum_{k=1}^{4}E_{k}\left(q_{k}({\bf m})-p_{k}\right)$; where $\{E_{k}\}$ are the non-degenerate increasing eigenvalues of $\mathbb{H}$, $\{p_{k}\}$ are the decreasing spectral values of $\rho_{\bf A}$, and $\{q_{k}({\bf m})\}$’s take values from the set $\{{\footnotesize\frac{1}{4}\left(1+(-1)^{i}{\bf m}^{\intercal}{\bf r}_{1}+(-1)^{j}\left|{\bf r}^{\intercal}_{2}+(-1)^{i}{\bf m}^{\intercal}{\bf T}\right|\right)}\}_{i,j=0}^{1}$ in decreasing order, and ${\bf m}$ is an arbitrary Bloch vector with $|{\bf m}|=1$.
Proof is provided in Appendix VI. There we also obtain a closed letter expression of this quantity for the states having completely mixed marginals which is stated in the following Corollary.
Corollary 1.
For a two-qubit state having completely mixed marginals, $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})=\omega(E_{1}+E_{2})/2+(1-\omega)(E_{3}+E_{4})/2-\sum_{k}p_{k}E_{k}$, where $2\omega:=1+\sqrt{\lambda}$, with $\lambda$ being the maximum eigenvalue of the matrix ${\bf TT}^{\intercal}$.
The plot depicted in Fig.1 exhibits the phenomena of energy locking in two-qubit Werner class of states. Notably, this phenomenon promises nontrivial usages of nonclassical correlation in devising efficient quantum batteries [55, 56, 57]. In multi-cell quantum batteries, stored energy can be lost due to various factors such as decoherence, global and local interactions, etc. For an isolated system the evolution is restricted to unitary processes only. Additionally, if interactions between individual systems are not allowed, energy drainage can occur only due to local interactions. Our findings indicate that energy storage is not possible in classically correlated states when local interactions and classical channels are available between pairs of isolated quantum batteries. In contrast, our Theorem 2 shows that for a system with a non-degenerate Hamiltonian, non-classical correlation in the form of discord can efficiently store energy in such a scenario.
Discussions.– We start by pointing out the recent work of Ref. [58]. There also the authors explored the role of nonclassical correlation in ergotropic work extraction. As established in their Theorem 1, the zero value of the ergotropic quantity implies the bipartite state to be classically correlated. But, the other way implication does not hold. Subsequently, this study is extended by considering generalised measurements [59], albeit the question of both way implication remains unresolved. In contrast, our Theorem 2 presents faithful quantifiers of nonclassical correlations for arbitrary bipartite systems. An interesting question is whether these quantities serve as valid measures of nonclassicality. Recall that, a correlation measure ($M$) is expected to fulfill the following three criteria [19]: (i) faithfulness, meaning $M(\rho_{\bf A})=0\Leftrightarrow\rho_{\bf A}$ is a CQ state, (ii) unitary invariance, indicating $M((u_{1}\otimes u_{2})(\rho_{\bf A})(u_{1}^{\dagger}\otimes u_{2}^{\dagger}))=M(\rho_{\bf A}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ u_{1}\in\mathrm{U}_{1}\leavevmode\nobreak\ \&\leavevmode\nobreak\ u_{2}\in\mathrm{U}_{2}$, and (iii) monotonicity, where $M((\mathcal{I}\otimes\Lambda)(\rho_{\bf A}))\leq M(\rho_{\bf A})$ holds for any CPTP map $\Lambda$ acting on the $A_{2}$ subsystem. While we have established that the first two properties are satisfied, the last one requires further study. Another intriguing question, which we leave for further exploration, is whether these nonclassical quantifiers are independent of the original quantum discord measure. Two quantifiers $M_{1}$ and $M_{2}$ are considered independent if there exist nonclassical states $\rho_{\bf A}$ and $\sigma_{\bf A}$ such that $M_{1}(\rho_{\bf A})>M_{1}(\sigma_{\bf A})$, while $M_{2}(\rho_{\bf A})<M_{2}(\sigma_{\bf A})$. Furthermore, although our current investigation primarily focuses on bipartite systems, the generality of our framework suggests potential for extension to multipartite systems.
Acknowledgements.
Acknowledgements: MA acknowledges the funding supported by the European Union (Project QURES- GA No. 101153001). SRC acknowledges support from University Grants Commission, India (reference no. 211610113404). MB acknowledges funding from the National Mission in Interdisciplinary Cyber-Physical systems from the Department of Science and Technology through the I-HUB Quantum Technology Foundation (Grant no: I-HUB/PDF/2021-22/008).